# Equation Discovery with Bayesian Spike-and-Slab Priors and Efficient Kernels

### v1

@@ 6 Empirical Results
For evaluation, we considered several benchmark ODEs/PDEs in the literature of data-driven equation discovery (Chen et al., 2021b, ; Sun et al.,, 2022). We examined whether KBASS can recover these equations from sparse, noisy measurements. We used the SE kernel and implemented KBASS with Jax (Frostig et al.,, 2018). We compared with state-of-the-art data-driven equation discovery methods, including SINDy (Brunton et al.,, 2016), PINN-SR (Chen et al., 2021b, ) and BSL (Sun et al.,, 2022). For the competing methods, we used the original implementation released by the authors. For a fair comparison, we carefully tuned every method to achieve the best performance (to our best effort). We leave the details about the settings and tuning of all the methods in Section B of Appendix.
Van del Pol (VDP) oscillator. We first tested the discovery of a nonlinear ODE system, $x_{t}=y,y_{t}=\mu(1-x^{2})y-x$, where $\mu=2.5$ and $x(0)=y(0)=1$. We used two test settings. The first setting employed only 10 evenly-spaced training examples in $[0,8]$ while the second setting used $25$ such examples. For each setting, we considered three noise levels: 0%, 1% and 20%. The dictionary includes the combination of the polynomials of each state variable up to 4th order, which gives $24$ candidates in total. To evaluate the discovery performance, we followed (Sun et al.,, 2022) to examine the normalized root-mean-square error (RMSE) of the weight estimation (including the relevant and irrelevant operators), and the precision and the recall of the selected operators.
@@
Figure 1: Solution and weight posterior estimation for the VDP equation with $10$ training examples (marked as green); (c) and (d) show the weight posterior for $x$ and $-x$, respectively. Their posterior selection probabilities were both estimated as $1.0$.
@@
Figure 4: Solution and weight posterior estimation on Allen-cahn equation; $w(\cdot)$ and $s(\cdot)$ denote the weight and selection indicator of the operator.
@@
where the initial condition is $u(x,0)=\exp(-(x-4)^{2})$. We tested with three viscosity values, $\nu=0.1$, $\nu=0.01$, and $\nu=0.005$. The dictionary consists of the polynomials and derivatives of $u$ up to 4th order and their combinations, which in total includes 24 candidate operators. For $\nu=0.1$, we first tested with 100 training examples on an evenly-spaced $10\times 10$ grid in the domain. We then tested with a denser dataset, including 400 examples at a $20\times 20$ evenly-spaced grid. Next, for $\nu=0.01$ and $\nu=0.005$, we used training examples at a $50\times 50$ evenly-spaced grid.
@@
$u_{t}=10^{-4}u_{xx}+5u-5u^{3},\;\;(x,t)\in[-1,1]\times[0,1],$ (18)
Real-world application. Finally, we ran our method on the data of a real-world predator-prey system. The dataset and experimental details are given in Section C of Appendix.
Results and discussion. We report the discovery performance of each method in Table 1, 2, 3, and 4 and Table 6 in Appendix. We can see that in all the settings, KBASS successfully recovered the equations, and the error of the operator weight estimate is small. As a comparison, SINDy failed to find the equations for every case. This might be due to that the sampled measurements (i.e., training examples) are too sparse. We found that SINDy started to successfully recover VDP, Lorenz 96, Burger’s ($\nu=0.1$), and KS equations when we increased to 50, 250, 1.6K and 384.5K examples respectively, which takes 5x and 21x, 16x and 240x of the sample size needed by KBASS. Nonetheless, the accuracy of the operator weight estimation is still much worse than KBASS; see the results in Table 7 of Appendix. PINN-SR failed to discover all the ODEs. This is consistent with the observation in (Sun et al.,, 2022). Though PINN-SR can exactly recover the Burger’s equation ($\nu=0.1$) with $10\times 10$ and $20\times 20$ examples, when adding a tiny amount of noise (1%), PINN-SR immediately failed; see Table 2(a). Similarly, BSL can exactly recover the Burger’s equation ($\nu=0.1$) using $20\times 20$ examples, with zero or 1% noise. However, it failed with 20% noise. Hence, it illustrates how KBASS is much more robust to data noise. In addition, both BSL and PINN-SR failed to find the Burger’s equation with $\nu=0.01$ and $\nu=0.005$, which are more challenging; see Table 2(b). So too did they for the KS and Allen-cahn equations; see Table 4. Note that although in (Chen et al., 2021b, ), PINN-SR successfully recovers a KS equation, it uses $320\times 101$ examples — 20x of the examples used by KBASS. We were not able to use PINN-SR to recover the KS equation tested in our experiment with the same amount of data. This might relate to our usage of a much larger domain, $[0,32\pi]\times[0,150]$ versus $[0,32\pi]\times[0,100]$.
Next, we showcased our solution prediction and posterior estimate of the operator weights in Fig. 1, 2, 3, 4, and Fig. 5, 6, 7 in Appendix. The shaded region in the solution prediction indicates the predictive standard deviation (scaled by a factor to be more clear). We used the Laplace’s approximation (see details in Section D of Appendix) to estimate the posterior of the solution prediction.
@@
We can see that the solution prediction by KBASS is quite accurate compared to the ground-truth, even the training data is very sparse and noisy, e.g., the prediction for Burger’s equation with $\nu=0.005$; see Fig 3. This might be (partly) due to that the accurate discovery of the equation in our E step can simultaneously guide the solution learning (M step). From the posterior estimate for the operator weights, we can see that the posterior mean is close to the ground-truth, while the Gaussian distribution allows us to further quantify the uncertainty. It is particularly interesting to see that in Fig. 3(b), KBASS successfully identifies the operator $u_{xx}$ that has a very small weight $10^{-4}$ and provides a Gaussian posterior approximation. Note that the posterior selection probability $q(s(u_{xx})=1)=1.0$, reflecting that KBASS has a very high confidence to select $u_{xx}$, even though the scale of the weight is very small. Hence, KBASS does not need to use a weight threshold for operator selection.
Computational efficiency. We examined the running time of each method. The results are given in Section E of Appendix. As shown in Table 8 of Appendix, KBASS takes much less time than BSL. For example, on Burger’s and KS, PINN-SR takes 9x and 20x running time but still failed to recover the KS equation. Together this shows our method also has a significant advantage in computational efficiency.
@@ 7 Conclusion
We have presented KBASS, an efficient kernel learning method with Bayesian spike-and-slab priors for data-driven equation discovery. Our method has higher sample efficiency and is more resistant to data noise than existing methods. KBASS has shown significant advantages on a series of PDE/ODE discovery tasks. Currently, KBASS is limited to a pre-specified operator dictionary. In the future, we plan to develop methods that can dynamically expand the dictionary to further improve the accuracy and robustness.
Appendix
Appendix A EP Spike-and-Slab Inference
A.1 EP Tutorial
@@
Because the exponential family is closed under multiplication and division, we can immediately obtain a closed-form approximate posterior $q(\bm{\theta})\approx p(\bm{\theta}|\mathcal{D})$ by merging the approximation terms in the RHS of (21), which is still a distribution in the exponential family.
@@ A.2 EP Method for Our Model
$p({\bf s},{\bf w},\mathcal{D})\propto\prod_{j=1}^{A}\text{Bern}(s_{j}|\rho_{0})\left({\bf s}_{j}\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-s_{j})\delta(w_{j})\right)\mathcal{N}({\bf h}|\bm{\Phi}{\bf w},\tau{\bf I}).$ (27)
@@
$s_{j}\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-s_{j})\delta(w_{j})\approx c_{j}\text{Bern}(s_{j}|\sigma(\rho_{j}))\mathcal{N}(w_{j}|\mu_{j},v_{j}).$ (28)
@@
$\widehat{\rho}_{j}=\sigma^{-1}(\rho_{0})+\rho_{j},\;\;\bm{\Sigma}=\left(\text{diag}^{-1}({\bf v})+\tau^{-1}\bm{\Phi}^{\top}\bm{\Phi}\right)^{-1},\bm{\beta}=\bm{\Sigma}\left(\frac{\bm{\mu}}{{\bf v}}+\tau^{-1}\bm{\Phi}^{\top}{\bf h}\right),$ (30)
@@
$\rho^{{\backslash}j}=\widehat{\rho}_{j}-\rho_{j},\left(v^{{\backslash}j}\right)^{-1}=[\bm{\Sigma}]_{jj}^{-1}-(v_{j})^{-1},\left(v^{{\backslash}j}\right)^{-1}\mu^{{\backslash}j}=\frac{\beta_{j}}{[\bm{\Sigma}]_{jj}}-\frac{\mu_{j}}{v_{j}}.$ (32)
@@
${\widetilde{p}}(s_{j},w_{j})$ $\propto q^{{\backslash}j}(s_{j},w_{j})\cdot\left(s_{j}\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-s_{j})\delta(w_{j})\right)$
@@
${\widetilde{p}}(s_{j})\propto\left(\sigma(\rho^{{\backslash}j})\mathcal{N}(\mu^{{\backslash}j}|0,v^{{\backslash}j}+\sigma_{0}^{2})\right)s_{j}+(1-\sigma(\rho^{{\backslash}j}))\left(\mathcal{N}(\mu^{{\backslash}j}|0,v^{{\backslash}j})\right)(1-s_{j}).$ (34)
@@
${\widetilde{p}}(w_{j})=\frac{1}{Z}(\sigma(\rho^{{\backslash}j})\mathcal{N}(w_{j}|\mu^{{\backslash}j},v^{{\backslash}j})\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-\sigma(\rho^{{\backslash}j}))\mathcal{N}(w_{j}|\mu^{{\backslash}j},v^{{\backslash}j})\delta(w_{j})).$ (37)
@@
Then, we obtain the global posterior,
@@ Appendix B Experimental Details
PINN-SR. We used the implementation (https://github.com/isds-neu/EQDiscovery) by the authors. PINN-SR first performs a pre-training of the NN from the training data, and then conducts alternating direction optimization (ADO) for joint equation discovery and solution learning. We tuned the number of NN layers from the range $[3,8]$, the NN width from range $[20,100]$, the number of iterations for ADO from range $[12,20]$, regularization strength from $[10^{-7},10^{-1}]$, L-BFGS pre-training iterations from $[10\text{K},80\text{K}]$, L-BFGS ADO training iterations from [1K, 2K], learning rate of ADAM from $[10^{-4},10^{-3}]$, and d_tol hyper-parameter from $[10^{-5},5]$. We varied the number of collocation points from $[10\text{K},160\text{K}]$.
@@
KBASS. We slightly tuned two hyper-parameters, the slab variance $\sigma_{0}^{2}$ from 0.5 to 5000 (very flat Gaussian slab), and the equation likelihood variance $\tau$ from $10^{-4}$ to $1$. The data likelihood variance $v$ in (9) (of the main paper) is obtained via cross-validation in the initial training (see the first step in Algorithm 1). The performance is not sensitive to other hyperparameters and so we fixed them. Specifically, the maximum number of EP-EM iteration was set to 200K and the tolerance level is $10^{-8}$. We also used relative change of natural parameters less than $10^{-5}$ as the condition to stop the EP inference at the E step. The solution estimation ${\mathcal{U}}$ was initialized as zero. We used an evenly-spaced mesh for each testing case, and the size is summarized in Table 5.
@@
Appendix C Real-world Predator-Prey Data
@@
which we used as the golden standard. We used the same setting as in (Sun et al.,, 2022) to test KBASS. The performance of the discovery is reported in Table 6. As we can see, KBASS not only correctly discovered the equation form, but also it achieves the smallest RMSE in the operator weight estimation. We also show the solution prediction of KBASS in Fig. 7.
@@
Appendix D Laplace’s approximation for Posterior Estimation

### v2

@@ 6 EMPIRICAL RESULTS
For evaluation, we considered several benchmark ODEs/PDEs in the literature of data-driven equation discovery (Chen et al., 2021b, ; Sun et al.,, 2022). We examined wether KBASS can recover these equations from sparse, noisy measurements. We used the SE kernel and implemented KBASS with Jax (Frostig et al.,, 2018). We compared with state-of-the-art data-driven equation discovery methods, including SINDy (Brunton et al.,, 2016), PINN-SR (Chen et al., 2021b, ) and BSL (Sun et al.,, 2022). For the competing methods, we used the original implementation released by the authors. For a fair comparison, we carefully tuned every method to achieve the best performance (to our best effort). We leave the details about the settings and tuning of all the methods in Appendix Section A.2.
Van del Pol (VDP) oscillator. We first tested the discovery of a nonlinear ODE system, $x_{t}=y,y_{t}=\mu(1-x^{2})y-x$ where $\mu=2.5$ and $x(0)=y(0)=1$. We used two test settings. The first setting employed only 10 evenly-spaced training examples in $[0,8]$ while the second setting used $25$ such examples. For each setting, we considered three noise levels: 0%, 1% and 20%. The dictionary includes the combination of the polynomials of each state variable up to 4th order, which gives $24$ candidates in total. To evaluate the discovery performance, we followed (Sun et al.,, 2022) to examine the normalized root-mean-square error (RMSE) of the weight estimation (including the relevant and irrelevant operators), and the precision and the recall of the selected operators.
@@
Figure 1: Solution and weight posterior estimation for the VDP equation with $10$ training examples (marked as green); (c) and (d) show the weight posterior for terms $y$ and $x$, respectively. Their posterior selection probabilities were both estimated as $1.0$.
@@
Figure 4: Solution and weight posterior estimation on Allen-cahn equation; $w(\cdot)$ and $s(\cdot)$ denote the weight and selection indicator of the operator. Note the ground-truth weight for $u_{xx}$ is $10^{-4}$.
@@
where the initial condition is $u(x,0)=\exp(-(x-4)^{2})$. We tested with three viscosity values, $\nu=0.1$, $\nu=0.01$, and $\nu=0.005$. The dictionary consists of the polynomials and derivatives of $u$ up to 4th order and their combinations, which in total includes 24 candidate operators. For $\nu=0.1$, we first tested with 100 training examples on an evenly-spaced $10\times 10$ grid in the domain. We then tested with a densor dataset, including 400 examples at a $20\times 20$ evenly-spaced grid. Next, for $\nu=0.01$ and $\nu=0.005$, we used training examples at a $50\times 50$ evenlly-spaced grid.
@@
$u_{t}=10^{-4}u_{xx}+5u-5u^{3},\;\;(x,t)\in[-1,1]\times[0,1],$
Real-world application. Finally, we ran our method on the data of a real-world predator-prey system. The dataset and experimental details are given in Appendix Section A.3.
Results and discussion. We report the discovery performance of each method in Table 1, 2, 3, and 4 and Appendix Table 6. We can see that in all the settings, KBASS successfully recovered the equations, and the error of the operator weight estimate is small. As a comparison, SINDy failed to find the equations for every case. This might be due to that the sampled measurements (i.e., training examples) are too sparse. We found that SINDy started to successfully recover VDP, Lorenz 96, Burger’s ($\nu=0.1$), and KS equations when we increased to 50, 250, 1.6K and 384.5K examples respectively, which takes 5x and 21x, 16x and 240x of the sample size needed by KBASS. Nonetheless, the accuracy of the operator weight estimation is still much worse than KBASS; see the results in Appendix Table 7. PINN-SR failed to discover all the ODEs. This is consistent with the observation in (Sun et al.,, 2022). Though PINN-SR can exactly recover the Burger’s equation ($\nu=0.1$) with $10\times 10$ and $20\times 20$ examples, when adding a tiny amount of noise (1%), PINN-SR immediately failed; see Table 3(a). Similarly, BSL can exactly recover the Burger’s equation ($\nu=0.1$) using $20\times 20$ examples, with zero or 1% noise. However, it failed with 20% noise. Hence, it illustrates that KBASS is much more robust to data noise. In addition, both BSL and PINN-SR failed to find the Burger’s equation with $\nu=0.01$ and $\nu=0.005$, which are more challenging; see Table 3(b). So too did they for the KS and Allen-cahn equations; see Table 4. Note that although in (Chen et al., 2021b, ), PINN-SR successfully recovers a KS equation, it uses $320\times 101$ examples — 20x of the examples used by KBASS. We were not able to use PINN-SR to recover the KS equation tested in our experiment even with the same amount of data. This might relate to our usage of a much larger domain, $[0,32\pi]\times[0,150]$ versus $[0,32\pi]\times[0,100]$.
Next, we showcased our solution prediction and posterior estimate of the operator weights in Fig. 1, 2, 3, 4, and Appendix Fig. 5, 6, 7. The shaded region in the solution prediction indicates the predictive standard deviation (scaled by a factor to be more clear). We used the Laplace’s approximation (see details in Appendix Section A.4) to estimate the posterior of the solution prediction.
We can see that the solution prediction by KBASS is quite accurate compared to the ground-truth, even when the training data is very sparse and noisy, e.g., the prediction for Burger’s equation with $\nu=0.005$; see Fig 3. This might be (partly) due to that the accurate discovery of the equation in our E step can simultaneously guide the solution learning (M step). From the posterior estimate for the operator weights, we can see that the posterior mean is close to the ground-truth, while the Gaussian distribution allows us to further quantify the uncertainty. It is particularly interesting to see that in Fig. 4(b), KBASS successfully identifies the operator $u_{xx}$ that has a very small weight $10^{-4}$ and provides a Gaussian posterior approximation. Note that the posterior selection probability $q(s(u_{xx})=1)=1.0$, reflecting that KBASS has a very high confidence to select $u_{xx}$, even though the scale of the weight is very small. Hence, KBASS does not need to use a weight threshold for operator selection.
Computational efficiency. We examined the running time of each method. The results are given in Appendix Section A.5. As shown in Appendix Table 8, KBASS takes much less time than BSL. For example, on Burger’s and KS , PINN-SR takes 9x and 20x running time but still failed to recover the KS equation. Together this shows our method also has a significant advantage in computational efficiency.
@@ 7 CONCLUSION
We have presented KBASS, an efficient kernel learning method with Bayesian spike-and-slab priors for data-driven equation discovery. Our method has higher sample efficiency and is more resistant to data noise than existing methods. KBASS has shown significant advantages on a series of PDE and ODE discovery tasks. Currently, KBASS is limited to a pre-specified operator dictionary. In the future, we plan to develop methods that can dynamically expand the dictionary to further improve the accuracy and robustness.
Acknowledgements
This work has been supported by MURI AFOSR grant FA9550-20-1-0358, NSF CAREER Award IIS-2046295 and NSF OAC-2311685.
Appendix A APPENDIX
A.1 EP SPIKE-AND-SLAB INFERENCE
A.1.1 EP Tutorial
@@
Because the exponential family is closed under multiplication and division, we can immediately obtain a closed-form approximate posterior $q(\bm{\theta})\approx p(\bm{\theta}|\mathcal{D})$ by merging the approximation terms in the R.H.S of (17), which is still a distribution in the exponential family.
@@ A.1.2 EP Method for Our Model
$p({\bf s},{\bf w},\mathcal{D})\propto\prod_{j=1}^{A}\text{Bern}(s_{j}|\rho_{0})\left(s_{j}\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-s_{j})\delta(w_{j})\right)\cdot\mathcal{N}({\bf h}|\bm{\Phi}{\bf w},\tau{\bf I}).$ (23)
@@
$s_{j}\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-s_{j})\delta(w_{j})\approx c_{j}\text{Bern}(s_{j}|\sigma(\rho_{j}))\mathcal{N}(w_{j}|\mu_{j},v_{j}),$ (24)
@@
$\widehat{\rho}_{j}=\sigma^{-1}(\rho_{0})+\rho_{j},\;\;\bm{\Sigma}=\left(\text{diag}^{-1}({\bf v})+\tau^{-1}\bm{\Phi}^{\top}\bm{\Phi}\right)^{-1},$
$\bm{\beta}=\bm{\Sigma}\left(\frac{\bm{\mu}}{{\bf v}}+\tau^{-1}\bm{\Phi}^{\top}{\bf h}\right),$ (26)
@@
$\rho^{{\backslash}j}$ $=\widehat{\rho}_{j}-\rho_{j},$
$\left(v^{{\backslash}j}\right)^{-1}$ $=[\bm{\Sigma}]_{jj}^{-1}-(v_{j})^{-1},$
$\left(v^{{\backslash}j}\right)^{-1}\mu^{{\backslash}j}$ $=\frac{\beta_{j}}{[\bm{\Sigma}]_{jj}}-\frac{\mu_{j}}{v_{j}}.$ (28)
@@
${\widetilde{p}}(s_{j},w_{j})$
$\propto q^{{\backslash}j}(s_{j},w_{j})\cdot\left(s_{j}\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-s_{j})\delta(w_{j})\right)$
@@
${\widetilde{p}}(s_{j})\propto\left(\sigma(\rho^{{\backslash}j})\mathcal{N}(\mu^{{\backslash}j}|0,v^{{\backslash}j}+\sigma_{0}^{2})\right)s_{j}+\left((1-\sigma(\rho^{{\backslash}j}))\mathcal{N}(\mu^{{\backslash}j}|0,v^{{\backslash}j})\right)(1-s_{j}).$ (30)
@@
${\widetilde{p}}(w_{j})=\frac{1}{Z}\left(\sigma(\rho^{{\backslash}j})\mathcal{N}(w_{j}|\mu^{{\backslash}j},v^{{\backslash}j})\mathcal{N}(w_{j}|0,\sigma^{2}_{0})+(1-\sigma(\rho^{{\backslash}j}))\mathcal{N}(w_{j}|\mu^{{\backslash}j},v^{{\backslash}j})\delta(w_{j})\right).$ (33)
@@
We obtain the global posterior approximation,
@@ A.2 EXPERIMENTAL DETAILS
PINN-SR. We used the implementation (https://github.com/isds-neu/EQDiscovery) of the authors. PINN-SR first performs a pre-training of the NN from the training data, and then conducts alternating direction optimization (ADO) for joint equation discovery and solution learning. We tuned the number of NN layers from the range $[3,8]$, the NN width from range $[20,100]$, the number of iterations for ADO from range $[12,20]$, regularization strength from $[10^{-7},10^{-1}]$, L-BFGS pre-training iterations from $[10\text{K},80\text{K}]$, L-BFGS ADO training iterations from [1K, 2K], learning rate of ADAM from $[10^{-4},10^{-3}]$, and d_tol hyper-parameter from $[10^{-5},5]$. We varied the number of collocation points from $[10\text{K},160\text{K}]$.
@@
KBASS. We tuned the slab variance $\sigma_{0}^{2}$ from 0.5 to 5000 (very flat Gaussian slab), and the equation likelihood variance $\tau$ from $10^{-4}$ to $1$. The data likelihood variance $v$ in (7) (of the main paper) is obtained via cross-validation in the initial training (see the first step in Algorithm 1). In the early iterations, we employed small values of $\alpha$, typically around $0.1$ or the smallest value of the current $q(s_{j}=1)$ among all $j$, to conservatively prune the operators. As the iteration progresses towards completion, we increased $\alpha$ to $0.5$ to conduct an unbiased pruning over the remaining operators, considering that our solution estimate is as optimal as possible. The performance is not sensitive to other hyperparameters and so we fixed them. Specifically, the maximum number of EP-EM iteration was set to 200K and the tolerance level is $10^{-8}$. We also used relative change of natural parameters less than $10^{-5}$ as the condition to stop the EP inference at the E step. The solution estimation ${\mathcal{U}}$ was initialized as zero. We used an evenly-spaced mesh for each testing case, and the size is summarized in Table 5.
@@
A.3 REAL-WORLD PREDATOR-PREY DATA
@@
which we used as the golden standard. We used the same setting as in (Sun et al.,, 2022) to test KBASS. The performance of the discovery is reported in Table 6. As we can see, KBASS not only correctly discovered the equation form, but also it achieves the smallest RMSE in the operator weight estimation. We also show the solution prediction of KBASS in Fig.  7.
@@
A.4 LAPLACE’S APPROXIMATION FOR POSTERIOR ESTIMATION